(D -2) x + Dy = 10sin2t

Dx + (D+2)y =0


1
Expert's answer
2021-06-07T17:16:32-0400

From


"(D-2)x +Dy =10 \sin t"

We have:


"Dx-2x+Dy=10 \sin t \quad \cdots (i)"


Also from


"Dx+(D+2)y =0"

We have:


"Dx+Dy+2y=0\\
Dy=-Dx-2y \quad \cdots (ii)"

Put (ii) in (i):

"Dx-2x-Dx-2y=10 \sin t\\
-2x-2y = 10 \sin t\\
-2(x+y)=10 \sin t\\
x+y=-5 \sin t\\"

Thus


"y=-x-5 \sin t"

is the solution